OpenStudy (r_josh10):

More 7th grade math questions wll an and medal

OpenStudy (r_josh10):

Here

OpenStudy (anonymous):

I = Prt

OpenStudy (anonymous):

I = interest P = principal r = rate t = time

OpenStudy (anonymous):

t = time in years

OpenStudy (r_josh10):

What next? @iGreen.

OpenStudy (anonymous):

Not sure

OpenStudy (anonymous):

Let me do a little research, hold on.

OpenStudy (r_josh10):

@Mindblast3r and @AnswerMyQuestions i need help

OpenStudy (anonymous):

Simple interest = (P*R*T)/100

OpenStudy (anonymous):

^^^^^^^^^^^

OpenStudy (anonymous):

I think we need to find r..

OpenStudy (anonymous):

No idea..

OpenStudy (r_josh10):

@iGreen. what grade are you in?

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

I never learned this though ;-;

OpenStudy (anonymous):

time (T) = 3 years Simple interest = 1575 Principal (P) = 5000. Plug into the formula and solve for R.

OpenStudy (anonymous):

Oh..that's right.

OpenStudy (anonymous):

I knew we had to do something like that..I was confused with the Simple Interest rate..anyway: I = Prt 1575 = (5000)(3)r Multiply 5000 * 3 @r_josh10

OpenStudy (r_josh10):

15000

OpenStudy (anonymous):

\[\frac{5000*R*3}{100} = 1575\] Solve for R

OpenStudy (anonymous):

Yes, now divide 15000 to both sides, 1575 / 15000 = ?

OpenStudy (anonymous):

Then multiply by 100.

OpenStudy (r_josh10):

10.5? @iGreen.

OpenStudy (anonymous):

multiply both sides by 100 first divide both sides by 15000 you get \[R= \frac{1575*100}{15000}\]

OpenStudy (anonymous):

absolutely correct

OpenStudy (anonymous):

Yep, you got it. 0.105 0.105 * 100 = 10.5%

OpenStudy (anonymous):

Gave both of ya medals ;)

OpenStudy (r_josh10):

Thank you guys i wish i could medal u both because you both helped me.

OpenStudy (anonymous):

It worked! :D

OpenStudy (r_josh10):

lol yes

OpenStudy (anonymous):

I gave one to you and @akitav too.